OpenStudy (anonymous):

find the solution set of each of the following linear equations. a) 7x -5y = 3 b) -8x1 + 2x2 - 5x3 + 6x4 = 1

OpenStudy (anonymous):

y=7/5 * x -3/5

OpenStudy (anonymous):

b) null

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

for a isnt it 3/7 + 5/7t y = t?

OpenStudy (anonymous):

true. also true. you have 2 variables and one equation therefor you can express one variable in terms of another.

OpenStudy (anonymous):

can you please show your work?

OpenStudy (anonymous):

7x -5y = 3 7x = 3+5y x = (3+5y)/7 => x=3/7+5/7 * y 7x -5y = 3 -5y = 3 - 7x y = (3-7x)/-5 y = -3/5 +7/5 * x

OpenStudy (anonymous):

good?

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

okk